What volume of methane at STP would be required to convert a 2.4 kg

kg block of ice at −16∘C to water at 16 ∘C, assuming that no heat is lost? The heat of fusion for ice is 6.01 kJmol−1. The heat capacity of ice is 2.108Jg−1∘C−1, the specific heat capacity of water is 4.184 Jg−1∘C−1 and the heat of combustion for methane is 891 kJmol−1.



1
Expert's answer
2021-02-04T07:04:38-0500

Q= mc∆T.

Q= 2400 x 6.01 x 16 = 230 784 J.

Q= 2400 x 2.108 x 16 = 80 947,2 J.

Q= 2400 x 4.184 x 16 = 160 665,6 J.

Q= 230784 + 80947.2 + 160665.6 = 472 396,8 J.

n= 472 396,8/891 x 1000 = 0,530187205 moles.

V= 0,530187205 x 22.4 = 11,8761934 L.